# A Note on Statistically Accurate Tabular Data Generation Using Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have shown promise in synthetic tabular data generation, yet existing methods struggle to preserve complex feature dependencies, particularly among categorical variables. This work introduces a probability-driven prompting approach that leverages LLMs to estimate conditional distributions, enabling more accurate and scalable data synthesis. The results highlight the potential of prompting probobility distributions to enhance the statistical fidelity of LLM-generated tabular data.
@@ 1 Introduction
In this study, we aim to enhance LLM-based synthetic tabular data generation by focusing on the accurate modeling of categorical variables. We propose a probability-driven prompting approach that estimates the distribution of categorical values and uses these estimates to generate synthetic data. This method seeks to preserve the statistical properties and interdependencies of the original data, thereby producing more realistic and useful synthetic datasets for various applications.
@@ 3 Results and Discussion
As can be seen in Fig. 1, the table-wide generation approach (panel b) broadly mimics the structure of the original data but exhibits slightly increased variance and over-smoothing across age groups, especially among underrepresented minorities. In contrast, the cell-by-cell method (panel c) fails to preserve population heterogeneity and demographic granularity, generating uniform values across all age groups and ethnicities, except for a subtle presence of ”Asian” and ”Black” categories.
To better reproduce the original data behavior, we propose a probability-driven prompting approach. It involves three key steps. First, the model receives a structured prompt that defines the dataset’s context, including relevant variables and categories (see Listing 3 in Appendix B). The model then predicts probability distributions for different categories based on the given context. Finally, these probabilities are used to randomly sample values for each row in a way that maintains realistic correlations between different attributes (see the pseudo-code in Listing 4). For example, in our California demographic data experiment, the new approach demonstrates its ability to capture age-dependent ethnicity distributions (see Fig.1d). By first generating age groups and then conditioning ethnicity probabilities on these age groups, the system produces synthetic data that better reflects the real-world correlation between age and ethnicity in California’s population, especially for dominant (”Latino” and ”White”) and minority groups, thus offering a superior balance between realism and diversity.
@@
This efficiency stands in sharp contrast to the cell-by-cell generation approach, where each record and each cell often requires a separate query, leading to substantial computational costs. While it is true that increasing the number of columns or introducing more granular categories can enlarge the distributional search space, thus potentially increasing the number of required prompts, this method still scales far more favorably. In such cases, although the number of required queries may asymptotically approach that of the cell-by-cell generation strategy, the cost remains bounded to distribution-level queries rather than record-level generation.
@@ 4 Summary
Furthermore, this approach exhibits a high degree of flexibility and adaptability. It can accommodate new categories or unseen feature values while maintaining logical consistency in the data - within the bounds of the model’s pre-training. If needed, the method can also be further fine-tuned on domain-specific datasets, enabling precise control over feature semantics and ensuring alignment with particular data generation objectives. Ultimately, this makes the method a interesting solution for scalable, accurate, and customizable synthetic tabular data generation.
@@
Age Group Race Original [17] This work Table-wide Cell-by-cell

### v2

@@ Abstract
Large language models (LLMs) have shown promise in synthetic tabular data generation, yet existing methods struggle to preserve complex feature dependencies, particularly among categorical variables. This work introduces a probability-driven prompting approach that leverages LLMs to estimate conditional distributions, enabling more accurate and scalable data synthesis. The results highlight the potential of prompting probability distributions to enhance the statistical fidelity of LLM-generated tabular data.
@@ 1 Introduction
In this study, our objective is to improve the generation of synthetic tabular data based on LLM by focusing on the accurate modeling of categorical variables. We propose a probability-driven prompting approach that estimates the distribution of categorical values and uses these estimates to generate synthetic data. This method seeks to preserve the statistical properties and interdependencies of the original data, producing more realistic and useful synthetic datasets for various applications.
@@ 3 Results and Discussion
As can be seen in Fig. 1, the table-wide generation approach (panel b) broadly mimics the structure of the original data, but exhibits slightly increased variance and over-smoothing across age groups, especially among underrepresented minorities. In contrast, the cell-by-cell method (panel c) fails to preserve population heterogeneity and demographic granularity, generating uniform values across all age and ethnicity groups, except for a subtle presence of ”Asian” and ”Black” categories.
To better reproduce the behavior of the original data, we propose a probability-driven prompting approach. It involves three key steps. First, the model receives a structured prompt that defines the dataset’s context, including relevant variables and categories (see Listing 3 in Appendix B). The model then predicts probability distributions for different categories based on the given context. Finally, these probabilities are used to randomly sample values for each row in a way that maintains realistic correlations between different attributes (see the pseudo-code in Listing 4). For example, in our California demographic data experiment, the new approach demonstrates its ability to capture age-dependent ethnicity distributions (see Fig.1d). By first generating age groups and then conditioning ethnicity probabilities on these age groups, the system produces synthetic data that better reflects the real-world correlation between age and ethnicity in California’s population, especially for dominant (”Latino” and ”White”) and minority groups, thus offering a superior balance between realism and diversity.
@@
This efficiency stands in sharp contrast to the cell-by-cell generation approach, where each record and each cell often require a separate query, leading to substantial computational costs. While it is true that increasing the number of columns or introducing more granular categories can enlarge the distributional search space, thus potentially increasing the number of required prompts, this method still scales far more favorably. In such cases, although the number of required queries may asymptotically approach that of the cell-by-cell generation strategy, the cost remains bounded to distribution-level queries rather than record-level generation.
@@ 4 Summary
Furthermore, this approach exhibits a high degree of flexibility and adaptability. It can accommodate new categories or unseen feature values while maintaining logical consistency in the data - within the bounds of the model’s pre-training. If needed, the method can also be further fine-tuned on domain-specific datasets, enabling precise control over feature semantics and ensuring alignment with particular data generation objectives. Ultimately, this makes the method an interesting solution for scalable, accurate, and customizable synthetic tabular data generation.
@@
Age Group Ethnicity Group Original [17] This work Table-wide Cell-by-cell